# End-to-end Streaming model for Low-Latency Speech Anonymization

### v1

@@ Abstract
keywords:
speaker anonymization, voice conversion, voice privacy, speech synthesis
@@ 1 Introduction
Existing machine learning (ML) based approaches to speaker anonymization follow a cascaded ASR-TTS architecture [1, 2]. An ASR module produces a text transcription that is speaker independent but eliminates emotional cues that may otherwise be of use for downstream applications. Moreover, existing systems for speaker anonymization are computationally heavy, operate in a non-streaming fashion, and/or have high latency on CPU devices. For speech anonymization to be used in the field, it must operate at real time (or faster), exhibit low latency, require minimal future context and be compatible with low-resource devices (e.g., smartphones).
To address these needs, we propose an end-to-end streaming model suitable for low-latency speaker anonymization. Our model draws inspiration from neural audio codecs [3, 4] for audio compression in low-resource streaming settings. Our proposed architecture consists of: (a) a streaming waveform encoder that generates a speaker-independent content representation from raw waveforms, (b) a pseudo-speaker generator that produces an anonymized speaker representation (or embedding) from the input speech, (c) a variance adapter that adds speaker, pitch and energy information to the content representation, and (d) a streaming decoder that consumes the output of the variance adapter and the corresponding speaker embedding to generate the final anonymized audio waveform.
Our system is trained as an autoencoder to reconstruct the input using pretrained speaker encoders [5, 6]. During inference, a pseudo-speaker generator produces a target speaker embedding with cosine distance greater than 0.3 from the source embedding, ensuring that the re-synthesized utterance sounds as if a different (i.e., anonymized) speaker had produced it. Additionally, the variance adapter is used to modulate pitch and energy values to further enhance privacy and control their similarity with the source audio. We show that our proposed system consisting of lightweight causal convolutional network can achieve similar performances as non-causal transformer networks which are computationally intensive.
We perform experiments on two versions of our model, a base version that can perform real-time streaming synthesis with a latency of 230ms and a lite version (having 0.1x the number of parameters) that further reduces latency to 66ms while maintaining state-of-the-art performance on naturalness, intelligibility, privacy and speaker identity transfer111 https://warisqr007.github.io/demos/stream-anonymization.
@@ 3 Method
The proposed system is illustrated in Figure 1. A reference waveform from the source speaker is passed to a pretrained speaker encoder to generate the source speaker embedding. The pseudo speaker generator receives the source speaker embedding and generates an anonymized speaker embedding. The content encoder receives streaming chunks of raw waveform and converts it into a hidden representation $z$ that contains the linguistic content disentangled from the speaker representation. The content information $z$ and the anonymized speaker embedding is passed to the variance adapter. The variance adapter injects pitch and energy values and then conditions the anonymized speaker embedding on the content representation $z$. The decoder receives the output of the variance adapter and the anonymized speaker embedding to generate the final anonymized waveform. We train two versions of our proposed system, a base and a lite version. Below, we describe each component of our system and the training procedure in detail.
@@
Figure 1: Block diagram of the proposed anonymization system. (a) training workflow (b) inference workflow.
3.1 Encoder
@@
Similar to [25], the content encoder predicts discrete speech units produced by discretizing the output speech representation from a pretrained HuBERT model [26] into one of N codewords or pseudo-labels. Our content encoder architecture follows that of HiFiGAN [27], except all transposed convolutions in HiFiGAN are replaced with strided causal convolutions to downsample the input waveform. Additionally, to support streaming applications, we replace all vanilla CNN layers in HiFiGAN with causal CNNs so that the prediction only considers the past context and does not rely on future audio frames. For both versions of our model (base and light), we use downsampling rates of $[2,2,4,4,5]$. The residual blocks have kernel sizes as $[3,7,11]$ with dilation rates as $[[[1,1],[3,1],[5,1]]*3]$ (please refer [27] for details). The difference between base and the lite version is the dimension of the hidden representation $z$ (the output of the encoder): 512 dimensions for the base version and 128 dimensions for the lite version.
3.2 Speaker embedding and speaker adapter
@@
Speaker verification or classification systems generally use speaker embeddings to represent the characteristics or timbre of a speaker’s voice. Widely used speaker encoders include the GE2E model [28], X-vectors [6] and ECAPA-TDNN [5]. Our systems concatenates embeddings generated from X-vectors and ECAPA-TDNN models, since these two models have been shown to complement each other [1].
We use a speaker adapter to condition the speaker embedding on the latent representations. The speaker adapter is based on adaptive instance normalization (adaIN) [29] and feature-wise linear modulation (FiLM) [30]. The conditioning goes as follows. First, we apply instance normalization to the input feature representation, and then transform it with scale and bias parameters learned through two 1D CNNs that take speaker embeddings as input.
3.3 Pseudo-speaker Generator
@@
To perform speaker anonymization, we use a pseudo-speaker generator that takes the original speaker embedding as input and outputs an artificially generated speaker embedding such that the generated anonymized speaker embedding has a cosine distance greater than 0.3 as compared with the original speaker embedding. Our pseudo-speaker generator follows the GAN-based framework proposed in [1] and is trained separately. The generator is trained to receive a random vector sampled from a standard normal distribution $N(0,1)$ as input and output a vector of the same shape as the original speaker embedding. The discriminator is trained to discriminate w.r.t the quadratic Wasserstein distance and transport cost [31] between the artificial and the original speaker embeddings.
3.4 Variance Adapter
@@
The variance adapter aims to add speaker, pitch, and energy (i.e., variance) information to the speaker-independent content representation and provides a way to control them [32]. The variance adapter consists of three modules: (a) speaker adapter, (b) pitch predictor, and (c) energy predictor. The speaker adapter conditions the speaker embedding on the content representation, and passes it to the pitch and energy predictors. During training, we use the ground-truth pitch and energy values to train the pitch and energy predictors. At inference time, we predict pitch and energy from the hidden embedding $z$. The predicted pitch and energy values are then conditioned on the $z$. The pitch and energy predictors have similar architecture consisting of a 2-layer 1D causal CNN with ReLU activation, followed by layer normalization and dropout layer and an additional 1D CNN (with kernel size 1) to project pitch and energy values on the content representation.
@@ 3.5 Decoder
The decoder follows the same design as HiFiGAN [27] and can be seen as a mirror-image of the content encoder with causal CNNs. Additionally, we condition the speaker embedding through the speaker adapter at the output of each residual block. For both versions of our model, we use upsampling rates of $[5,4,4,2,2]$. The residual blocks have kernel sizes as $[3,7,11]$ with dilation rates as $[[[1,1],[3,1],[5,1]]*3]$.
@@ 3.6 Training
The content encoder is trained to predict the pseudo-labels at the output of the HuBERT module using cross-entropy loss. We prevent gradient flow (i.e., back-propagation) from the decoder to the encoder, to ensure that the speaker information is not leaked via the content representation. The pitch and energy predictor in the variance adapter apply mean-squared error loss for pitch and energy prediction. At the output of the decoder, Following the HiFiGAN architecture [27], we apply a combination of adversarial losses at the output of the decoder, including feature loss, multi-period discriminator loss, multi-scale discriminator loss, multi-resolution STFT loss [33] and Mel-spectrogram reconstruction loss. These discriminators have similar architecture as those in HiFiGAN, including similar weighting schemes to compute the total loss.
@@ 4 Experimental setup
We trained our system on the LibriTTS corpus [34] following guidelines for the Voice Privacy Challenge 2022 (VPC22) [35]. All our experimental results are presented on the LibriTTS dev and test set, which were not part of the training. We use a pretrained HuBERT [26] [^2] https://github.com/facebookresearch/fairseq and extract the the output from its 9th layer. We set the number of cluster centroids to 200. For all our experiments, we use a sampling rate of 16 kHz and batch size of 16 with the AdamW optimizer with a learning rate of $2*10^{-4}$ annealed down to $10^{-5}$ by exponential scheduling. The encoder is first pretrained for 300k steps (for training stability), and then trained together with the decoder for an additional 800k steps. The pretrained speaker encoders were taken from speechbrain [36]. The pseudo speaker embedding generator follows the training procedure described in [1]. All our models are trained using two NVIDIA Tesla V100 GPUs.
@@ 5.1 Synthesis Latency
Our pretrained HuBERT model produces speech frames at 50Hz, so the smallest chunk size that our model can process is 20 ms. In this section, we present the synthesis latency for the base and lite versions for our model, for various chunk sizes between 20ms and 140ms on both CPU and GPU devices. We define latency as the sum of chunk size and the average time taken by the model to synthesize the corresponding chunk. For a system to be real-time, the latency should be less than twice the chunk size. Results are summarized in Table 1. On GPU, our base version can operate in real-time for the chunk size of 40ms with a latency of 64ms, while on CPU the base model can be real-time for chunk size of 120ms with a latency of 230ms. In case of the lite version, the model is real-time for chunk size of 20ms with a latency of 38ms on GPU and can operate in real-time for 40ms with latency of 66ms on the CPU. For our next set of experiments we set the chunk size of 120ms and 40ms for the base and lite versions respectively.
Table 1: Synthesis latency on CPU and GPU
@@ 5.2 Synthesis Quality
We use DNSMOS [40] as an objective measure of naturalness for our experiments. DNSMOS provides three ratings for quality of speech (SIG), noise (BAK), and overall (OVRL). Additionally, we assess the intelligibility of synthesized speech through Word Error Rate (WER). Table 2, summarizes results for the five baselines and the proposed systems. In terms of DNSMOS, our models achieve similar ratings as Diff-VC, QuickVC, and B3 across the three measures, and comparable or better than the source speech. In terms of intelligibility, our systems achieve comparable WER to that of B3, and superior to the rest, even though our models operate in a causal fashion with far more limited context.
We verified the synthesis quality of our two models through listening tests on Amazon Mechanical Turk (AMT). Namely, participants (N=20) were asked to rate the acoustic quality of utterances using a standard 5-point scale mean opinion score (MOS) (5: excellent, 1: bad). Each listener rated utterances synthesized using the base and lite models, as well as original utterances (20 for each). Results are shown in Table 3. Both systems (base and lite) obtained comparable ratings of MOS as the original utterances. We see a difference of 0.1 between the MOS score of base and lite versions, but we didn’t find them to be significant ($\ll$ 0.001). It is noteworthy that while the lite version has 0.1x number of parameters, it achieves nearly the same synthesis quality as the base version.
@@ 5.3 Speaker Anonymization
To assess speaker-anonymization performance, we report Equal Error Rate (EER) on the speaker verification model (ASV) in the VoicePrivacy 2024 Challenge github (see section 4). ASV tests are conducted for the following two scenarios, (a) ignorant, where we only anonymize the trial data (O-A), or (b) lazy-informed, where we anonymize both enrollment and trial data but use different target speakers (A-A). Results are shown in Table 4. For both the ignorant and lazy-informed scenario, our models achieves similar performance as B3 and outperforms VPC22 B2. Although our base model performs slightly worse than B3, the differences are not significantly different between them ($\ll$ 0.001).
To corroborate these results, we conducted an AB listening test on AMT. Participants were presented with two audio samples, one from a speaker in the enrollment set, and the second sample from one of three options: (a) a different utterance from the same speaker from the trial set, (b) an utterance from a different speaker from the trial set, or (c) another utterance of the same speaker from the trial set but anonymized through our lite version of the system. Then, participants had to decide if both samples were from the same speaker, and rate the confidence in their decision using a 7-point scale (7: extremely confident; 5: quite a bit confident; 3: somewhat confident; 1: not confident at all). Each listener rated 20 AB pairs per scenario. Results are summarized in Table 5. In settings (a) and (b), listeners could easily identify whether the recording were from the same or different speakers (81 % ) with high confidence (5.71). In setting (c) However, the anonymized trial data obtained obtained similar rating as in (b), indicating that proposed system was able to anonymize the trial recordings.

### v2

@@ Abstract
Index Terms—  speaker anonymization, voice conversion, voice privacy, speech synthesis
@@ 1 Introduction
Existing machine learning (ML) based approaches to speaker anonymization follow a cascaded automatic speech recognition (ASR) – text-to-speech (TTS) architecture [1, 2]. An ASR module produces a text transcription that is speaker independent but eliminates emotional cues that may otherwise be of use for downstream applications. Moreover, existing systems for speaker anonymization are computationally heavy, operate in a non-streaming fashion, and/or have high latency on CPU devices as opposed to GPUs. For speech anonymization to be used in the field, it must operate at real time (or faster), exhibit low latency, require minimal future context and be compatible with low-resource devices (e.g., smartphones).
To address these needs, we propose an end-to-end streaming model suitable for low-latency speaker anonymization. Our model draws inspiration from neural audio codecs [3, 4] for audio compression in low-resource streaming settings. Our key strategy that enables streaming is to replace traditional non-causal computationally intensive networks (e.g., ASR or self-supervised learning based models) for encoding linguistic content with a lightweight convolutional neural network (CNN) based architecture. Our proposed architecture consists of: (a) a streaming waveform encoder that generates a speaker-independent content representation from waveforms, (b) a pseudo-speaker generator that produces an anonymized speaker representation (i.e., an embedding) from the input speech, (c) a speaker/variance adapter that adds speaker, pitch and energy information to the content representation, and (d) a streaming decoder that consumes the speaker/variance adapted linguistic representation and the corresponding speaker embedding to generate the final anonymized audio waveform. Our system is trained in an auto-encoder fashion, which reconstructs the input conditioned on the speaker embeddings generated using pre-trained speaker encoders [5, 6]. During inference, a pseudo-speaker generator produces a target speaker embedding with cosine distance greater than 0.3 from the source embedding, ensuring that the re-synthesized utterance sounds as if a different (i.e., anonymized) speaker had produced it. Additionally, the speaker/variance adapter is used to modulate pitch and energy values to further enhance privacy and control the similarity of the synthesized speech with the source audio. We show that our lightweight convolutional neural network (CNN) based architecture achieves similar performance as traditional content encoders.
We perform experiments on two versions of our model, a Base version that can perform real-time streaming synthesis with a latency of $230ms$ and a Lite version (having 0.1x the number of parameters) that further reduces latency to $66ms$ while maintaining state-of-the-art performance on naturalness, intelligibility, privacy and speaker identity transfer [^1] https://warisqr007.github.io/demos/stream-anonymization/.
@@ 3 Method
The proposed system is illustrated in Figure 1. Anonymization takes place in two steps, (1) generating a fixed (i.e., off-line) anonymized speaker embedding personalized to the source speaker and, (2) using this fixed anonymized speaker embedding and the streaming speech synthesizer to synthesize anonymized speech that only preserves the linguistic content of the source speech. To generate the anonymized embedding, a reference waveform from the source speaker is passed to the pre-trained speaker encoder, which then produces the source speaker embedding. The pseudo speaker generator receives this source speaker embedding and generates the anonymized version (see Figure 1b). To synthesise speech signals, the content encoder receives streaming chunks of waveform and converts it into a hidden representation $z$ that contains the linguistic content disentangled from the speaker representation. The content information $z$ and the anonymized speaker embedding (generated in the previous step) is passed to the speaker/variance adapter. The speaker/variance adapter, first, conditions the anonymized speaker embedding on the content representation $z$ and then adds pitch and energy values. The decoder receives the output of the speaker/variance adapter and the anonymized speaker embedding to generate the final anonymized waveform.
We train two versions of our proposed system, a base and a lite version. Below, we describe each component of our system and the training procedure in detail.
@@
Fig. 1: Block diagram of the proposed anonymization system. (a) training workflow (b) inference workflow (c) speaker/variance adapter.
3.1 Content encoder
@@
The content encoder consumes the wav signal to predict discrete speech units produced by discretizing the output speech representation from a pretrained HuBERT model [25] into one of N codewords or pseudo-labels [26]. Our content encoder architecture follows that of HiFiGAN [27], except all transposed convolutions in HiFiGAN are replaced with strided causal convolutions to downsample the input waveform. Additionally, to support streaming applications, we replace all vanilla CNN layers in HiFiGAN with causal CNNs so that the prediction only considers the past context and does not rely on future audio frames. For both versions of our model (base and lite), we use downsampling rates of $[2,2,4,4,5]$. The residual blocks have kernel sizes as $[3,7,11]$ with dilation rates as $[[[1,1],[3,1],[5,1]]*3]$ (please refer [27] for details). The difference between base and the lite version is the dimension of the hidden representation $z$ (the output of the encoder): 512 dimensions for the base version and 128 dimensions for the lite version.
3.2 Speaker encoder and pseudo-speaker generator
@@
Speaker verification or classification systems generally use speaker embeddings to represent the characteristics or timbre of a speaker’s voice. Widely used speaker encoders include the GE2E model [28], X-vectors [6] and ECAPA-TDNN [5]. Our system concatenates embeddings generated from X-vectors and ECAPA-TDNN models, since these two models have been shown to complement each other [1].
@@
To perform speaker anonymization, we use a pseudo-speaker generator that takes the original speaker embedding as input and outputs an artificially generated speaker embedding such that the generated anonymized speaker embedding has a cosine distance greater than 0.3 as compared with the original speaker embedding. Our pseudo-speaker generator follows a GAN-based architecture [1] and is trained separately. The generator is trained to receive a random vector sampled from a standard normal distribution $N(0,1)$ as input and output a vector of the same shape as the original speaker embedding. The discriminator is trained to discriminate w.r.t the quadratic Wasserstein distance and transport cost [29] between the artificial and the original speaker embeddings.
3.3 Speaker/Variance adapter
The speaker/variance adapter aims to add speaker, pitch, and energy (i.e., variance) information to the speaker-independent content representation and provides a way to control them [30]. The speaker/variance adapter consists of three modules: (a) speaker adapter, (b) pitch predictor, and (c) energy predictor (see Figure 1c).
The speaker adapter conditions the speaker embedding on the content representation $z$, and passes it to the pitch and energy predictors. The speaker adapter is based on adaptive instance normalization (adaIN) [31] and feature-wise linear modulation (FiLM) [32]. The conditioning goes as follows. First, we apply instance normalization to the input feature representation, and then transform it with scale and bias parameters learned through two 1D CNNs that take speaker embeddings as input. The use of instance normalization was motivated by a prior work [10] that showed instance normalization being helpful in removing residual speaker information.
@@ 3.3 Speaker/Variance adapter
Pitch and energy predictor estimate pitch and energy values based on speaker adapted content representation $z$. During training, we use the ground-truth pitch and energy values to train the pitch and energy predictors. At inference, the output of the pitch and energy predictor are added to the speaker adapted $z$. The pitch and energy predictors have similar architecture consisting of a 2-layered 1D causal CNNs (kernal size 3) with ReLU activation, followed by layer normalization and dropout layer and an additional 1D CNN (with kernel size 1) to project pitch and energy values on the latent representation.
@@ 3.4 Decoder
The decoder follows the same design and training procedure as HiFiGAN [27] and can be seen as a mirror-image of the content encoder. Similar to the content encoder, all vanilla CNNs are replaced with causal CNNs. The decoder receives speaker/variance adapted latent representation along with speaker embedding and directly generates waveform signal without any intermediate mel-spectrogram generation. We additionally adapt the output of each residual block of the decoder to the speaker embedding. In our experiments, we observed that doing so gave better speaker transfer performance. For both versions of $base$ and $lite$ version of our model, we use upsampling rates of $[5,4,4,2,2]$. The residual blocks have kernel sizes as $[3,7,11]$ with dilation rates as $[[[1,1],[3,1],[5,1]]*3]$.
@@ 3.5 Training
The training workflow is described in Figure 1a. The proposed system is trained end-to-end similar to an autoencoder, reconstucting the same waveform at output that fed as input. The content encoder is trained to predict the pseudo-labels generated through a HuBERT/Kmean module using cross-entropy loss. The pitch and energy predictor in the variance adapter apply mean-squared error loss for pitch and energy prediction. At the output of the decoder, Following the HiFiGAN architecture [27], we apply a combination of adversarial losses at the output of the decoder, including feature loss, multi-period discriminator loss, multi-scale discriminator loss, multi-resolution STFT loss [33] and mel-spectrogram reconstruction loss. These discriminators have similar architecture as those in HiFiGAN, including weighting schemes to compute the total decoder loss. The final training loss is the summation of content loss, pitch/energy error loss and decoder losses. We apply a stop-gradient operation to prevent gradient flow (i.e., back-propagation) from the decoder to the encoder, to ensure that the speaker information is not leaked via the content representation. This operation effectively decouples it from the rest of the system; in other words, the encoder and the rest of the system can potentially be trained sequentially as two independent modules.
@@ 4 Experimental setup
We trained our system on the LibriTTS corpus [34] following guidelines for the Voice Privacy Challenge 2022 (VPC22) [35]. All our experimental results are presented on the LibriTTS dev and test set, which were not part of the training. We use a pretrained HuBERT-base [^2] https://github.com/facebookresearch/fairseq and extract the the output from its 9th layer. We set the number of cluster centroids to 200. For all our experiments, we use a sampling rate of 16 kHz and batch size of 16 with the AdamW optimizer with a learning rate of $2*10^{-4}$ annealed down to $10^{-5}$ by exponential scheduling. The encoder is first pretrained for 300k steps (for training stability), and then trained together with the decoder for an additional 800k steps. The pretrained speaker encoders were taken from speechbrain [36]. The pseudo speaker embedding generator follows the training procedure described in [1] and trained on VoxCeleb 1 and 2 [37, 38]. All our models are trained using two NVIDIA Tesla V100 GPUs for approximately two weeks.
@@ 5.1 Synthesis Latency
Our pretrained HuBERT model produces speech frames at $50Hz$, so the smallest chunk size that our model can process is $20ms$. In this section, we present the synthesis latency and the real-time factors (RTF) for the base and lite versions for our model, for various chunk sizes between $20ms$ and $140ms$ on both CPU and GPU devices. Latency is defined as the sum of the chunk size and the average time the model takes to synthesize that chunk. RTF is the ratio of the system’s average processing time to the chunk size. For a system to be real-time, the latency should be less than twice the chunk size, meaning RTF should be less than 1. Results are summarized in Table 1. On GPU, our base version can operate in real-time for the chunk size of $40ms$ with a latency of $64ms$, while on CPU the base model can be real-time for chunk size of $120ms$ with a latency of $230ms$. In case of the lite version, the model is real-time for chunk size of $20ms$ with a latency of $38ms$ on GPU and can operate in real-time for $40ms$ with latency of $66ms$ on the CPU. For our test set of experiments we set the chunk size of $120ms$ and $40ms$ for the base and lite versions respectively.
Table 1: Synthesis latency and real-time factors (RTF) on CPU and GPU devices.
@@ 5.2 Synthesis Quality
We use DNSMOS [42] as an objective measure of naturalness for our experiments. DNSMOS provides three ratings for quality of speech (SIG), noise (BAK), and overall (OVRL). Additionally, we assess the intelligibility of synthesized speech through Word Error Rate (WER). We calculated WER using an ASR consisting of a CRDNN based acoustic model and a transformer-based language model that uses CTC and attention decoders. Table 2, summarizes results for the five baselines and the proposed systems. In terms of DNSMOS, our models achieve comparable ratings as Diff-VC, QuickVC, and B3 across the three measures, and comparable or better than the source speech. In terms of intelligibility, our systems achieve comparable WER to that of B3, and superior to the rest, even though our models operate in a causal fashion with far more limited context.
We verified the synthesis quality of our two models through listening tests on Amazon Mechanical Turk (AMT). Namely, participants (N=20) were asked to rate the acoustic quality of utterances using a standard 5-point scale mean opinion score (MOS) [rating, speech quality, level of distortion]: [5, excellent, imperceptible] — [4, good, just perceptible but not annoying] — [3, fair, perceptible but slightly annoying] — [2, poor, annoying but not objectionable ] — [1, bad, very annoying and objectionable]. Each listener rated utterances synthesized using the base and lite models, as well as original utterances (20 for each). Results are shown in Table 3. Both systems (base and lite) obtained comparable ratings of MOS as the original utterances. We see a difference of 0.1 between the MOS score of base and lite versions, but we didn’t find them to be significant. It is noteworthy that while the lite version has 0.1x number of parameters, it achieves nearly the same synthesis quality as the base version.
@@ 5.3 Speaker Anonymization
To assess speaker-anonymization performance, we report Equal Error Rate (EER) on the speaker verification model (ASV) in the VoicePrivacy 2024 Challenge github (see section 4). ASV tests are conducted for the following two scenarios, (a) ignorant, where we only anonymize the trial data (O-A), or (b) lazy-informed, where we anonymize both enrollment and trial data but use different target speakers (A-A). Results are shown in Table 4. For both the ignorant and lazy-informed scenario, our models achieves similar performance as B3 and outperforms VPC22 B2. Although our base model performs slightly worse than B3, the differences are not significantly different between them ($p=0.13$).
To corroborate these results, we conducted an AB listening test on AMT. Participants were presented with two audio samples, one from a speaker in the enrollment set, and the second sample from one of three options: (a) a different utterance from the same speaker from the trial set, (b) an utterance from a different speaker from the trial set, or (c) another utterance of the same speaker from the trial set but anonymized through our lite version of the system. Then, participants had to decide if both samples were from the same speaker, and rate the confidence in their decision using a 7-point scale (7: extremely confident; 5: quite a bit confident; 3: somewhat confident; 1: not confident at all). Each listener rated 20 AB pairs per scenario. Results are summarized in Table 5. In settings (a) and (b), listeners could easily identify whether the recording were from the same or different speakers (81 % ) with high confidence (5.71). In setting (c), however, the anonymized trial data obtained obtained similar rating as in (b), indicating that proposed system was able to anonymize the trial recordings.